News and Updates

News and Updates

About the blog

Read below our regular monthly / quarterly updates of work in progress and operations of the diesel locomotives based at the Bodmin and Wenford Railway.

Read more about the loco's and the group on our full website by clicking here.

May - June 2017

UpdatesPosted by Administrator Sun, July 02, 2017 14:14:14

Loco Status

47306 ‘The Sapper’ - In overhaul.

D3452 – Stored awaiting overhaul.

P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.

Operational dates this period.

13th May English Electric Special TT Day with 37142 & 50042

21st May x2 return runs Loco 37142 Yellow TT.

9th June Private charter 50042

17th & 18th June Heritage Transport Festival 37142.

24th June Mixed Traffic Day, x8 return trips shared diesel 37142 & steam.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Drivers - During this period, two additional drivers have passed out on our class 50 and one more on our class 37 locomotives, which should mean greater flexibility to run 50042 in particular.

Future running dates.

22nd July 37142 Lilac TT x6 return trips.

5th Aug TBC Lilac TT x6 return trips.

17th Aug TBC Pub Quiz evening.

19th Aug TBC Lilac TT x6 return trips.

2nd Sept TBC Lilac TT x6 return trips.

Maintenance for Service Duties

33110

33110 has seen little use during this period and is B&WR standby loco.

37142

Following the successful (temporary) repair to the coolant header tank, 37142 has once again settled into giving a steady reliable performance with no issues to report, passing fit to run exams.

50042 ‘Triumph’

50042 'Triumph' awaits the right away with the Private charter clays 9/6/17

Triumph was due to take part in the heritage transport weekend, however during the private charter a developing fault occurred with the train brake leaking beneath Cab 2 floor. Taken out of service to investigate, found the pipe was about to completely fail in the usual EE way through the cab floor meaning 50042 had to be withdrawn from the heritage weekend until repairs could be made.

With the cab floor already removed, two volunteers spent the best part of one hot June Saturday tackling the pipe fittings in one of the worst locations. Persistence, skill and at times a large hammer won through with the repaired pipe work being tested in the late afternoon.

Our regular exams highlighted a new issue in that the coupling pins were starting to wear and bend, so both have now been replaced.

We hope to run 50042 during one or two of our approaching summer turns.

In Overhaul

47306 ‘The Sapper’

The big news with ‘The Sapper’ is the generator is back and installed!


Early on Friday the 23rd June MacSalvors lifted the 7.5 ton generator from the lorry that had returned it to Bodmin and swung it into 47306 once more. By 10am the lift was all over and the rest of the day was spent bolting the generator down and coupling it to the engine. Rotation deflection tolerances were then checked and are well with limits.


The push now is to rebuild everything that was removed for the generator removal, namely the exhaust system, turbo, silencer, air intake and coolant pipework. Then we should hear 47306 start up once more.


Meanwhile the Radiator louvers have been completed and are back in position, Cab 2 internally is progressing with strip back and priming, along with more external bodywork areas also moving forwards.









  • Comments(0)//blog.bodmin-heritage-diesels.co.uk/#post29

March - April 2017

UpdatesPosted by Administrator Mon, May 01, 2017 09:00:15

Loco Status

47306 ‘The Sapper’ - In overhaul.

D3452 – Stored awaiting overhaul.

P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.

Operational dates this period.

5th March Driver Experience Course 37142.

18th March x5 return runs 37142 top & tail with steam Red TT.

19th March x5 return runs 37142 top & tail with steam Red TT.

9th April x2 return runs Loco 37142 Yellow TT.

22nd April Driver Experience Course 50042.

23rd April x2 return runs Loco 33110 Yellow TT.

29th April x6 return runs Loco 37142 Lilac TT.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Future running dates.

13th May English Electric Special TT Day with 37142 & 50042

21st May x2 return runs Loco 37142 or 33110 Yellow TT.

17th & 18th June Heritage Transport Festival.

24th June Mixed Traffic Day, x8 return trips shared diesel & steam.

Maintenance for Service Duties

33110

No issues this period, used on ash disposal runs, station shunts and stand in for 37142 on the 23rd April service see below.


37142

We start March with 37142 booked for a driver experience course, which ran successfully and recruited a new member to the diesel volunteers, welcome David.

A & B exams were carried out in early March with 37142 passed fit to run.

The usual March half price weekend saw 37142 in use both days top & tailing with our steam colleagues at the start of the 2017 operating season.

Post the yellow Sunday timetable on the 9th April we became aware that the water level in the coolant header tank was starting to drop by a small amount. Whilst topping the level up prior to service on the 23rd April, the small leak became a hole in the bottom of the tank very quickly!

Resulting in 37142 declared failed and 33110 being used instead for the Sunday duties. Inspections luckily found a way to carryout repairs (temporary) without the need to remove the engine roof, disconnect pipework and remove the tank etc. Hopefully keeping 37142 in service for some time to come, while the group look to effecting an overhaul or replacement of this tank.

29th April saw the repaired and passed fit to run 37142, hauling our first of the year Lilac timetables.

50042 ‘Triumph’

On the 18th April following its winter hibernation, 50042 was extracted from Parkway shed by 37142, prepped and fired up, running up fine and passing a fit to run exam. Hauling 37142 back up to Bodmin General its first haulage since the Xmas installation of the new KV10 main generator controller and last Octobers service train.

A booked driver experience course was undertaken on the 22nd April with three delighted participants and ‘Triumph’ certainly lived up to its name, driving the best it ever has in preservation from one of the group members who recalls how it arrived at Bodmin back in 1992!

Some minor modifications have been undertaken to the temperature controller circuits that should mean a longer life for the smaller relays involved in controlling the bigger 110v dc contactors.

A new fault has developed that means the exhauster speed up circuit for Cab 1 (Downhill at Bodmin) does not work, initial investigation looks like the wire that runs full length of the loco has failed, meaning at present drivers have to be careful not to over brake going down our gradients.

In Overhaul

47306 ‘The Sapper’

Cab 1 externally has been receiving attention with the marker plate removed from that end too, but this time welding in inserts to make good rather than a full replacement like cab 2. The process of filling & sanding back of this cab outside body work has also commenced, revealing the outline of its pre-TOPS numbers D1767!

Meanwhile Cab 2 exterior is making great progress with a first coat of gloss now on most surfaces and work inside the cab of stripping back and preparing has also recommenced now the engine room is nearly fully painted out.

Overhaul work on ‘A’ side rad louvers frames x2 continues and should see them reinstated before too long. As does the servicing of the numerous batteries and associated battery links.

Our contractor has been on & off fabricating new parts of the engine room roof supports and mating surfaces that had been cut away as too badly corroded to recover. Along with reports that the contractor ‘up’ country is not far away from finishing the generator overhaul, which may well be heading back this way soon.

Engine rocker covers are being cleaned, painted and new gaskets fitted as part of the process of working our way through the ‘to do’ list. The removed Turbo has been getting some servicing and fresh paint too while it’s easy to get at being off the loco. The failed internal lights timer has had the failed component identified, replaced and tested ready for re-installation later on.

All in all a very busy couple of months!

Next month English Electric Day 13th May, both 37142 & 50042 running an intensive timetable with hopefully 08444 available for 'Driver for a £5'









  • Comments(0)//blog.bodmin-heritage-diesels.co.uk/#post28

Jan - Feb 2017

UpdatesPosted by Administrator Sun, February 26, 2017 13:30:13

Loco Status

47306 ‘The Sapper’ - In overhaul.

D3452 – Stored awaiting overhaul.

P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.

Operational dates this period.

None this period.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Future running dates.

18th March x5 return runs 37142 top & tail with steam Red TT.

19th March x5 return runs 37142 top & tail with steam Red TT.

9th April x2 return runs Loco TBD Yellow TT.

23rd April x2 return runs Loco TBD Yellow TT.

29th April x6 return runs Loco TBD Lilac TT.

Maintenance for Service Duties

33110

Used for a couple of minor shunts and 50042 test run see below.

37142

Over Xmas 37142 saw a trip to the engine shed pit for bogie, under chassis, traction motor inspections, cleaning and servicing. Further work has seen an ‘A’ & ‘B’ Exam completed and the two leaking axle box gaskets renewed along with a grease purge ready for the 2017 season start.

With its radiator fan cover in place over the winter, 37142 has stayed relatively dry and fired up Mid-February on a cold day with flakes of snow falling. No issues at all found on a test run to Bodmin Parkway and back.

Photo James Goodhall, 37142 warming up from a 'Cold' start

50042 ‘Triumph’

Triumph was not forgotten over Xmas with a trip down to Parkway shed using 33110, saw us exchange out the old KV10 main generator power control module and fit the newly constructed one.

Whilst two members worked on this, another two carried out parts of an ‘A’ & ‘B’ Exam plus some servicing. Coupling up 33110 to 50042, the latter was extracted from the shed, prepared and started with testing of the new KV10 carried out. Declared working, 50042 was given a work out with 33110 acting as a load, whilst more monitoring of 50042 power handling was undertaken on our 1:40 gradient. The new KV10 replacement works well and early signs are that 50042 is handling power better than it has done since preservation with us at B&WR.

Photo 50042 New KV10 module installed and first checks being carried out on it.

In Overhaul

47306 ‘The Sapper’

There is light at the end of the tunnel as we are again putting things back on 47306 rather than taking them off! This last couple of months has seen a lot of effort going into the Radiator fan grills which open and close slats by hydraulic oil pressure. All four frames have been lowered down off the top of the loco. The first two B side have been stripped down, cleaned back to bare metal, primed and painted, including the many slats. Reassembled and manually tested prior to being installed back over the top of the radiators. The second set of frames and slats ‘A side’ are already progressing and receiving the same treatment.

With the generator out, the whole area underneath has been cleaned, painted and new rubber insulation laid around the cable junction area. Further floor panels have been cut, welded and painted around the engine area completing the floor repairs in the engine room.

Work also continues on repairing exhaust system parts as well as making the new engine room roof strips that the roof sits on and seals.

Cab1 work continues on the wetter days and the head code panel has also now been removed from this end too which like Cab 2 is badly corroded so will probably be replaced.

The news from the contractors that took receipt of the generator set, is outside of the banding starting to come away, the generator is in relatively good condition. It will still undergo a cleaning process, minor repairs, insulation varnishing, baking before being reassembled and returned.

There is of course lots of sanding, filling and painting along with a host of other minor repair work going on. Volunteer man hours of work on 47306 now exceed 3,500 since start of overhaul Nov 2016!













  • Comments(0)//blog.bodmin-heritage-diesels.co.uk/#post27

Oct - Dec 2016

UpdatesPosted by Administrator Mon, December 19, 2016 20:57:17

Loco Status

47306 ‘The Sapper’ - In overhaul.

D3452 – Stored awaiting overhaul.

P403D ‘Denise’, 08444, 33110, 37142 & 50042 ‘Triumph’ – In Traffic.

Operational dates this period.

23rd Oct x3 return runs 37142 Yellow TT.

29th Oct x5 return runs 37142 top & tail with steam Red TT.

30th Oct x5 return runs 50042 top & tail with steam Red TT.

2 x DECs (Driver Experience Courses) one each for 37142 & 50042.

Various dates – 08444, D3452 & 33110 - Permanent way and station shunt duties.

Future running dates.

No further service runs planned for 2016.

Awaiting release of 2017 timetable.

Maintenance for Service Duties

33110

Monitoring of the oil temperature continues but with 33110 now on light load / use with winter work the condition is not seen. Adjustments have been made to the hydraulic temperature valve that controls the radiator fan speed, have also improved the issue.

A second battery cell has failed causing 33110 to only just turn over on start cranking. The cell has been exchanged out with a spare one and 33110 now cranks over correctly.

37142

Following its last turn end of October, 37142 has had a cover placed over the Radiator fan grill to help prevent rain water ingress over the winter. Battery levels checked and topped up plus a topping up charge ready for the winter layup. It is planned to do some servicing work over Xmas and New Year ready for 2017 season.

50042 ‘Triumph’

Likewise post its last turn on the half price weekend at the end of Oct, 50042 was light engine run down to Parkway shed and is now stored dry.

Some tweaking of the radiator shutters has been done to aid the warm up period to be shorter hopefully on cooler days.

Over the Xmas break we are hoping to install the newly built and bench tested KV10 module, if installation is successful then this new module will be commissioned and tested to control the main generator. Once again thank you to those who donated money for this new module which should see 50042 more reliable in the future. A second KV10 module (Yes there is two inside) is also planned to be made and installed when funds allow, further donations to the group most welcome!

P403D ‘Denise’

Is settling in as yard shunter at Bodmin General.

D3452

The class 10 shunter is now awaiting a slot for overhaul. During November the shunter was moved to Bodmin Parkway shed for dry storage by 08444 making it the first time in over ten years that D3452 has been down the Parkway line and away from Bodmin General.

08444

Continues largely in service as yard shunter & Permanent Way duties.

In Overhaul

47306 ‘The Sapper’

It’s now a year since 47306 started its first major overhaul with the Bodmin Mainline Diesel Group, it’s amazing to see what has already been achieved in that time, nearly all by volunteer efforts.

In the last blog update we mentioned we were preparing for the generator to be removed. This occurred in November after many phone calls by the owner negotiating the right lorry and crane-age at the right price for its lift and road trip ‘Up North’. We had to manufacture our own armature support bracket to allow the generator and engine to be separated, a sturdy piece of iron work on its own! The lift itself went well and the weather even gave us a weather window on a very wet and windy day. Meaning the 7.5 ton generator is now away with the contractor being assessed for overhaul.

With the generator removed, working inside the engine room has become a lot easier! Cleaning and scrapping back the generator bed pan has been completed along with painting it out. Most of the internal sides in this area have also been cleaned, primed, undercoated and glossed.

Work continues on Cab 1 inside and out with the inside starting to see the finish post. A lot of filling and sanding has gone into Cab 2 and A side externally, with pink primer for the aluminium cab along with half the A side now in a first coat of anti-corrosion green primer. Work has also been undertaken on Cab 2 roof including making up new aerial base parts that had rotted out.

Work also continues on cleaning and servicing parts of the engine removed so far, rocker covers are all being gradually cycled through a clean-up and repainting. The Turbo is receiving attention as is the parts of the air intake & exhaust system removed, including welding up cracks and facing off the surfaces ready for new exhaust seals.

The Diesel group will be gathering on a couple of days over the Xmas / New Year break, if anyone is thinking of lending us a hand, why not make contact (contact@bodmin-heritage-diesels.co.uk) with us and see when we are going to be at Bodmin General.



  • Comments(0)//blog.bodmin-heritage-diesels.co.uk/#post26
Next »